Find a number which is 150 greater than its opposite. And check.
Answer by stanbon(75887)

Find a number which is 150 greater than its opposite. And check.
Let the number be "x"; the opposite is "-x".
EQUATION:
x-(-x) = 150
2x = 150
x = 75
-x = -75
-----------
Checking:
75-(-75) = 150
150 = 150
Cheers,
Stan H.